Shift report 20030127 16:00-24:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Mon Jan 27 2003 - 23:52:02 EST

  • Next message: Apache: "Shift report 20030128 00:00-08:00"
    ------------------------------------------------------------
    Date: Mon Jan 27 23:52:02 2003, Shift: 16:00-24:00
    Supervisor: PS, with: 
    Log book#13 pages 1 - 2
    
    Shift Summary:
    ------------------------------------------------------------
    1. 4h of beam on this shift, namely from 16:00 to 20:23
    Beam was dumped and RHIC people worked on deuteron ramp before the next store.
    
    2. Data runs: 
       MRS 90deg., 6B and FS 3deg. 1/2A: 7662, 7663, 7665 (588k trig#6 for this settings)
       MRS   same         FS 3deg. 1/4A: 7667, 7668 (417 trig#6).
    
      Other runs: 
      7671 - pedestal run
      7672 - TPC pedestal run,
      7675 - DVM run.
    
    3. No detector problem.
      
    ------------------------------------------------------------
    


    This archive was generated by hypermail 2.1.5 : Mon Jan 27 2003 - 23:52:41 EST